OpenStudy (anonymous):

Explain the difference between the solutions to 2x+3y=4, 2x+3y>4, 2x+3y>4.

OpenStudy (anonymous):

the 2 and 3 r the same equation

ganeshie8 (ganeshie8):

you mean them to be : 2x+3y=4, 2x+3y < 4, 2x+3y>4 ?

OpenStudy (anonymous):

2x+3y=4 2x+3y < 4 2x+3y>4 And to describe the difference between their solutions. That's what the question gives me!

ganeshie8 (ganeshie8):

Okay, you must be knowing that 2x+3y=4 is a straight line

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

all the points on that line satisfy that equation 2x+3y=4. so all the points on that line are solutions to that first equation

ganeshie8 (ganeshie8):

2x+3y < 4 since its a < symbol, intuitively enough, solutions for this wud fall below the line

ganeshie8 (ganeshie8):